OpenStudy (anonymous):

I am Stuck. Help. MEDAL!!! What is the length of stack H D with bar on top when BH = 20 in., BC = 12 in., and DC = 14 in.? Round your answer to the nearest tenth of an inch. in.

OpenStudy (anonymous):

Image

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@Loser66 can you help me.

OpenStudy (xapproachesinfinity):

just the use of Pythatgorean theorem couple of times

OpenStudy (loser66):

@xapproachesinfinity is a perfect guy for this

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

ok let come from a clearer picture! you recognize HDC is right triangle yes

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

we are given DC but we need HC to do the so called Pythagoras yes

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

now what we do is look at the triangle HCB it is also a right triangle and we are given BC and BH so we can find HC

OpenStudy (anonymous):

ok how?

OpenStudy (xapproachesinfinity):

so then we have the following:\ HC^2+BC^2=BH^2 HC^2=BH^2-BC^2 HC^2=20^2-12^2

OpenStudy (xapproachesinfinity):

now tell me what is after you take the square root

OpenStudy (xapproachesinfinity):

you have a calculator with right

OpenStudy (xapproachesinfinity):

i got HC=16

OpenStudy (anonymous):

me too

OpenStudy (xapproachesinfinity):

now we go back to our first triangle HDC now we have HC=16 and DC=14 we use again that same theorem to find HD HD^2+DC^2=HC^2 HD^2=HC^2-DC^2 HD^2=16^2-14^2

OpenStudy (xapproachesinfinity):

let see what you got

OpenStudy (anonymous):

60?

OpenStudy (xapproachesinfinity):

sqrt of 60 not just 60 because i told you to take square root after that too

OpenStudy (xapproachesinfinity):

so then HD=7.75 inch

OpenStudy (anonymous):

is that the answer?

OpenStudy (xapproachesinfinity):

if we wanted the decimal, well in a way we want the decimal since we measuring length

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

can you help me with three more?

OpenStudy (xapproachesinfinity):

let see post in a different section though

OpenStudy (anonymous):

ok I will tag you.

OpenStudy (xapproachesinfinity):

ok